# FiniteTimeBibliography.bib

### v1

@@ 1.2.2 Bypassing energy levels with finite-time autocorrelators
To motivate this relation, let us first consider a system with nondegenerate energy levels $E_{n}$, and some observable of interest $A$, whose thermal value we again initially assume to be constant throughout the spectrum (i.e., “global thermalization”). The connected autocorrelator of $\hat{A}$, which is given by the autocorrelator of $\widehat{\delta A}=(\hat{A}-A_{\mathrm{th}}\hat{\mathds{1}})$, can be averaged over infinite time to give
@@
This is related to the Mazur-Suzuki equality [MazurPhysica, SuzukiPhysica] for finite-dimensional quantum systems, which has also been implicitly used for ETH in e.g. [VenutiLiu, KawamotoStrongETH]. In particular, the (approximate) vanishing of the time-averaged autocorrelator (up to some normalization that we will presently ignore) implies eigenstate thermalization in the sense of Eq. (3), and the reverse implication also holds. From our point of view, however, the infinite time average is a major obstacle due to its inaccessibility (for reasons similar to those connecting ETH to thermalization, see the discussion around Eq. (6)). Quantitatively, for the $T\to\infty$ limit to begin approaching the right hand side, we need
@@
Once again, in the previously mentioned mathematical literature on semiclassical chaos [Zelditch, Sunada], the connection obtained between classical ergodicity and eigenstate thermalization in quantized classical systems proceeds through classical autocorrelators as in Eq. (1). As these are insensitive to the quantization of the spectrum, the existence of these results again strongly suggests that accessible, finite-time autocorrelators should be able to constrain thermalization in fully quantum systems.
@@
Our solution, which generalizes a more specific technique in the semiclassical literature [Zelditch], is to consider time-averages of autocorrelators weighted by non-negative functions $w_{+}(t)\geq 0$ whose Fourier transforms $\widetilde{w}_{+}(\delta E)\geq 0$ are also non-negative [^4] The semiclassical approach, e.g. [Zelditch], is to integrate $\operatorname{Tr}[\widehat{\delta A}(t_{1})\widehat{\delta A}(t_{2})]$ over $t_{1},t_{2}\in[-T,T]$, which is contained as a special case of our approach, and then take the classical limit $\hbar\to 0$, and then $T\to\infty$.. For convenience, we call these functions “completely positive” for want of better terminology (which is quite distinct from completely positive maps on density operators [NielsenChuang]). As an aside, such functions also play an important role in the fast scrambling problem, and more generally in the question of obtaining many-body speed limits related to the energy-time uncertainty principle for different classes of systems [dynamicalqspeedlimit, dynamicalqfastscrambling].
@@
If we consider completely positive time averages, we obtain in place of the traditional autocorrelators in Eq. (14):
@@
Here, all terms on the right hand side are non-negative, so a decay of the autocorrelator must imply the smallness of nonvanishing terms. If $w_{+}(t)$ has support on a range of times $t\in[-T_{w},T_{w}]$, then we show that $\hat{A}$ satisfies energy-band thermalization over all corresponding inverse energy scales (Proposition 5.1 in Sec. 5):
$[\hat{A}]_{\Delta E}\approx A_{\mathrm{th}}\hat{\mathds{1}},\text{ for all }\Delta E\ll\frac{2\pi}{T_{w}},$ (16)
which, by Eq. (9), also implies that time-averaged thermalization is achieved for almost all basis states (Corollary 5.2),
@@
In most of the text, we restrict ourselves to autocorrelators of projectors $\hat{\Pi}_{A}$ as the “simplest” observables which can be most directly accessed in experiments via projective measurements [NielsenChuang], noting that other more complicated observables can be expanded in terms of projectors to their eigenbasis (e.g. $\hat{A}=\sum_{a}A_{a}\hat{\Pi}_{a}$) and are usually inferred from these projective measurements. For projectors, it suffices to prepare a single initial state $\hat{\rho}_{A}\propto\hat{\Pi}_{A}$ to determine their thermalization in (almost) all possible initial states. A key takeaway from this approach is then that the dynamics of a projector observable in a single quantum state fully determines whether it thermalizes in arbitrary “physical” initial states.
@@
Second, we illustrate in App. B how our results may be used to show the thermalization of local observables in almost all basis states for any choice of basis in dual-unitary quantum circuits, in which autocorrelators can be computed exactly [ProsenErgodic, ClaeysErgodicCircuits, ArulCircuit], with thermal values that are uniform across all states and do not depend on any conserved quantities such as energy.
@@ 1.2.3 Interference effects for energy shells and conserved charges
Further, the conventional theoretical approach of regularizing, e.g., with a thermal density operator to focus on a narrow energy window is unsuitable for our purposes, because it would require showing properties of this external thermal state that are unlikely to be accessible (even for a rigorously justified regularization strategy in Ref. [dynamicalqfastscrambling]). To illustrate the problem, let us say that $\hat{A}=\hat{a}_{1}\otimes\hat{\mathds{1}}_{(N-1)}$ is a single-particle observable $\hat{a}_{1}$ on particle $1$. One strategy for restricting to some narrow range of energies, often useful in rigorous statements on thermalization [dynamicalqfastscrambling, ETHNecessity], is to initialize the remaining $(N-1)$ particles in some state $\hat{\rho}\propto\hat{\mathds{1}}_{1}\otimes\hat{\rho}_{N-1}$ with narrow energy support and measure the autocorrelator of $\widehat{\delta a}_{1}=\hat{a}_{1}-a_{\mathrm{th}}\hat{\mathds{1}}_{1}$ in this state. In attempting to derive expressions analogous to Eq. (15), we get:
@@
where $A_{\mathrm{th}}(\mathcal{E})$ is the thermal value in an energy shell of range $\mathcal{E}$ (assumed constant within the energy shell). Then we specifically show that a class of interference measures, which we generically call “quantum dynamical echoes”, related to Loschmidt echoes [LoschmidtEcho1, LoschmidtEcho2, LoschmidtEcho3],
@@
can be time-averaged with completely positive weight functions $w_{+}(t)$ and $v_{+}(t)$ to constrain energy-band thermalization within the energy shell spanning $\mathcal{E}$, with any desired energy-band width $\Delta E$. This rigorously connects the properties of the energy eigenstates to certain Loschmidt echoes, which has so far remained one of the missing links between two key objects of interest that fall under the umbrella of “quantum chaos” [Haake], which we take to refer to a collection of loosely related notions with clear physical distinctions that are often used to study complex quantum systems (see, e.g., [webofquantumchaos] for a summary of some links). We also show that the microscopic properties of energy levels may once again be bypassed completely, by directly relating the decay of these echoes to the thermalization of any physical basis states within the energy shell $\mathcal{E}$, as described for time-averaged thermalization in Summary 6.3, with its generalization to thermal equilibrium in Sec. 7.
@@
This strategy can also account for macroscopic conserved charges [^5] As all Hamiltonians conserve linear combinations of the energy projectors $\lvert E_{n}\rangle\langle E_{n}\rvert$ (including all projectors in degenerate eigenspaces), by a conserved charge we strictly just mean some observable $\hat{Q}$ formed by such a linear combination — which commutes with the Hamiltonian $[\hat{H},\hat{Q}]=0$ — on which the thermal value of our observable $\hat{A}$ of interest may strongly depend. By “macroscopic”, we mean a conserved charge with an inaccessibly large number of eigenvalues. For discrete conserved charges (with an accessible number of eigenvalues), it should be straightforward to project onto a desired eigenspace of the charge.: in the presence of a single conserved charge $\hat{Q}$ (for example) such that the thermal value $A_{\mathrm{th}}(\mathcal{E},\mathcal{Q})$ depends [DAlessio2016, NonabelianETH] on $Q$, being effectively constant over a range of eigenvalues $Q\in\mathcal{Q}$, we can update Eq. (20) to the form:
@@
Here, it is not essential for the charge $\hat{Q}_{\text{exp}}$ that may be accessible in experiments to be the exact conserved charge $\hat{Q}$, but just for the transformation $\exp(-i\hat{Q}_{\text{exp}}s)$ generated by it for accessible values of $s$ to be sufficiently close to the true symmetry transformation $\exp(-i\hat{Q}s)$ to allow a determination of the above echo to within experimental errors.
In summary, we develop a conceptual framework for quantum statistical mechanics that relies entirely on the measurable dynamical properties of a single state (per observable), as illustrated through explicit measurement protocols. This parallels what we find to be the simplest dynamical justification for classical statistical mechanics [KhinchinStatMech] (Eq. (1)), but with additional interesting consequences of interference effects in the quantum case that allow us to explicitly address finite time intervals, a finite energy resolution, and the presence of accessible conserved charges in an exact manner.
@@ 1.3 Organization of this paper
Sec. 3 reviews the conventional quantum statement of eigenstate thermalization with infinite time averages, but also our version (Definition 3.2) of a stronger statement for thermalization in degenerate eigenspaces that seems not to be widely known at all, whose variants have appeared before in Refs. [Sunada, VenutiLiu, TumulkaEsT]. Despite its relevance only for infinite timescales, we will use this notion to motivate energy-band thermalization by regarding energy bands as “approximate” degenerate eigenspaces over small timescales that are not sensitive to their resolution.
@@
For the Appendices, Appendix A derives a finite-time Mazur-Suzuki inequality for a finite dimensional quantum system, rigorously bounding completely positive time averages of autocorrelators in terms of approximately conserved charges. Appendix B illustrates an application of our results to show the thermalization of local observables in almost all (physical) initial states of dual-unitary quantum circuits. Appendix C analyzes the connection between energy-band thermalization and eigenspace thermalization, and concludes that eigenspace thermalization is not always accessible even given energy-band thermalization; consequently, attempting to bypass the energy levels and eigenstates entirely may be in our best interest for an experimentally accessible formulation of quantum statistical mechanics. Finally, as all our proofs are relatively straightforward (and repeated) applications of the triangle and Cauchy-Schwarz inequalities [ByronFuller], they are relegated to Appendix D, although what we perceive to be nontrivial physical statements are derived or explained in the course of the main text or the relevant Appendices.
@@ 2 Classical “eigenstate” thermalization without ergodicity
Now, with a primary view of motivating our quantum results, we will describe one possible approach to the classical autocorrelator, Eq. (1), as the determiner of thermalization in classical dynamical systems. Rather than take a more direct route [KhinchinStatMech], we will follow a somewhat contrived path by first considering a classical notion of eigenstate thermalization as a possible explanation for thermalization. While classical eigenstate thermalization has been discussed before in, e.g., Ref. [VenutiLiu] assuming the ergodic hypothesis, it is crucial for our purposes to formulate this notion for more general non-ergodic systems to resemble the quantum situation and demonstrate independence from ergodicity.
@@ 2.1 Setup: classical ergodicity
Figure 2: Schematic depiction of thermalization without ergodicity in a classical phase space $\mathcal{P}$ with $5$ ergodic subsets $\mathcal{P}_{k}$. The observable $\Pi_{A}$ is considered to be equally distributed over these ergodic subsets (i.e., shows eigenstate thermalization), and must therefore thermalize in almost all initial states. The observable $\Pi_{B}$ is restricted to a few of these subsets, and cannot thermalize almost everywhere. Separately, this figure also provides schematic intuition for the Mazur-Suzuki inequality [Eq. (18) and App. A], where the conserved quantities are regarded to be projectors onto the ergodic subspaces $Q_{k}=\Pi_{\mathcal{P}_{k}}$, among which we have access to, say, $k=3,4,5$. Here, $\Pi_{B}$ has a large overlap with these conserved quantities and its autocorrelator may fail to decay to its thermal value by the inequality, while $\Pi_{A}$ has sufficiently low overlap with all these conserved quantities to still allow autocorrelator thermalization (which in turn would imply eigenstate thermalization as well as thermalization).
Ergodicity is a difficult property to either prove or verify in any sufficiently large phase space, not least because of the sensitivity required to ensure that every infinitesimal region of $\mathcal{P}$ is visited by generic trajectories (in a $3$D gas of $N\to\infty$ particles, this would require having precise knowledge of the $3N$ positions and momenta of every particle). Instead of relying on ergodicity, let us ask when Eq. (29) can be true for a given measurement $A$ in a non-ergodic system. This is of extreme physical relevance: all nontrivial Hamiltonian systems are non-ergodic due to the conservation of energy, as well as the frequent presence of various other conservation laws (such as total momentum or angular momentum, depending on the system under consideration and nature of interactions). Further, experimental uncertainties ensure that any initial state is statistically distributed over a range of energies $\Delta E$. Moreover, as ergodicity in the dynamical sense is difficult (in most cases, near-impossible) to establish for a sufficiently complex system, so a given system may as well be nonergodic for all practical purposes. In this case, the phase space can be decomposed into an unknown number of subsets $\mathcal{P}_{k}$:
@@
each of which is ergodic. These are typically surfaces of fixed energy, as well as fixed values of other conserved quantities (if present). Through a slight misuse of terminology that anticipates quantum mechanics, we will call the sets $\mathcal{P}_{k}$ — or more precisely, uniform distributions supported entirely on one of these sets — the “energy eigenstates” of the classical Hamiltonian (for example, because they correspond to the classical limit of the quantum energy eigenstates, and satisfy similar properties such as having a definite value of energy and being invariant under time evolution [^6] This parallel can be taken further in the Koopman-von Neumann framework [HalmosErgodic] of classical Hilbert spaces comprised of functions on $\mathcal{P}$, in which classical dynamics has well-defined eigenstates.). An identification between constant energy surfaces and eigenstates was proposed in Ref. [VenutiLiu] in a related context, assuming ergodicity on each energy surface; for our perspective, it is instead crucial that we identify eigenstates with the subsets $\mathcal{P}_{k}$ especially if the system is non-ergodic.
For simplicity, we will assume that there is a finite number $M$ of $\mathcal{P}_{k}$, each with nonzero measure [^7] Any subsets of zero measure can be trivially absorbed into any nonzero measure subset without affecting its ergodicity. $\mu(\mathcal{P}_{k})>0$; formally, this simplifies the technical details of several arguments, and the case of a continuum of energy surfaces (typical in classical systems) can be recovered via a continuum limit $M\to\infty$ without altering any of our conclusions. This assumption allows us to focus on the physics of the problem, rather than the formal definition of induced measures on subsets and their integration. It also further anticipates quantum mechanics, in which there is a minimum phase space volume [StechelMeasure, StechelHeller] connected to the inverse purity of a pure state. With this simplification, due to the ergodicity of each subset, any observable $\Pi_{A}$ satisfies
@@
If Eq. (34) is satisfied, we say that the observable $\Pi_{A}$ shows eigenstate thermalization: its expectation value in every ergodic subset $\mathcal{P}_{k}$ is the thermal value $\mu(A)/\mu(\mathcal{P})$. Though classically such a notion is not particularly useful, because it tends to be the case that a typical accessible observable $A$ attains its thermal value in almost all states $x\in\mathcal{P})$ in the phase space (and therefore trivially for the energy eigenstates) [KhinchinStatMech, GallavottiErgodic], we will continue to formally entertain this notion in anticipation of later sections.
@@
See App. D.1.2. The intuition behind the proof (which is reminiscent of a proof of semiclassical wavefunction ergodicity theorems [ZelditchTransition, Sunada, Zelditch], which we will revisit for quantum thermalization) is as follows. The difference between the left and right hand sides of Eq. (35),

### v2

@@ 1.2.2 Bypassing energy levels with finite-time autocorrelators
[…]
@@
This is a version of the Mazur-Suzuki equality [48, 49] for finite-dimensional quantum systems, which has also been implicitly used for ETH in e.g. [50, 51]. In particular, the (approximate) vanishing of the time-averaged autocorrelator (up to some normalization that we will presently ignore) implies eigenstate thermalization in the sense of Eq. (2), and the reverse implication also holds. From our point of view, however, the infinite time average is a major obstacle due to its inaccessibility (for reasons similar to those connecting ETH to thermalization, see the discussion around Eq. (6)). Quantitatively, for the $T\to\infty$ limit to begin approaching the right hand side, we need
@@
Once again, in the previously mentioned mathematical literature on semiclassical chaos [46, 41], the connection obtained between classical ergodicity and eigenstate thermalization in quantized classical systems proceeds through classical autocorrelators as in Eq. (1). As these are insensitive to the quantization of the spectrum, the existence of these results again strongly suggests that accessible, finite-time autocorrelators should be able to constrain thermalization in fully quantum systems.
@@
Our solution, which generalizes a more specific technique in the semiclassical literature [46], is to consider time-averages of autocorrelators weighted by non-negative functions $w_{+}(t)\geq 0$ whose Fourier transforms $\widetilde{w}_{+}(\delta E)\geq 0$ are also non-negative [^4] The semiclassical approach, e.g. [46], is to integrate $\operatorname{Tr}[\widehat{\delta A}(t_{1})\widehat{\delta A}(t_{2})]$ over $t_{1},t_{2}\in[-T,T]$, which is contained as a special case of our approach, and then take the classical limit $\hbar\to 0$, and then $T\to\infty$.. For convenience, we call these functions “completely positive” for want of better terminology (which is quite distinct from completely positive maps on density operators [52]). As an aside, such functions also play an important role in the fast scrambling problem, and more generally in the question of obtaining many-body speed limits related to the energy-time uncertainty principle for different classes of systems [27, 28].
@@
If we consider completely positive time averages, we obtain, in place of the traditional autocorrelators in Eq. (14), the weighted autocorrelator:
@@
Here, all terms on the right hand side are non-negative, so a decay of the autocorrelator must imply the smallness of nonvanishing terms. If the autocorrelator decays over a timescale $T_{w}$, i.e., when $w_{+}(t)$ has support on a range of times $t\in[-T_{w},T_{w}]$, then we show that $\hat{A}$ satisfies energy-band thermalization over all corresponding inverse energy scales (Proposition 5.1 in Sec. 5):
[…]
which, by Eq. (9), also implies that time-averaged thermalization is necessarily achieved for almost all basis states over longer but comparable timescales to the decay of the autocorrelator (Corollary 5.2),
@@
In most of the text, we restrict ourselves to autocorrelators of projectors $\hat{\Pi}_{A}$ as the “simplest” observables which can be most directly accessed in experiments via projective measurements [52], noting that other more complicated observables can be expanded in terms of projectors to their eigenbasis (e.g. $\hat{A}=\sum_{a}A_{a}\hat{\Pi}_{a}$) and are usually inferred from these projective measurements. For projectors, it suffices to prepare a single initial state $\hat{\rho}_{A}\propto\hat{\Pi}_{A}$ to determine their thermalization in (almost) all possible initial states. A key takeaway from this approach is then that the dynamics of a projector observable in a single quantum state fully determines whether it thermalizes in arbitrary “physical” initial states.
@@
Second, we illustrate in App. B how our results may be used to show the thermalization of local observables in almost all basis states for any choice of basis in dual-unitary quantum circuits, in which autocorrelators can be computed exactly [55, 56, 57], with thermal values that are uniform across all states and do not depend on any conserved quantities such as energy.
@@ 1.2.3 Interference effects for energy shells and conserved charges
Further, the conventional theoretical approach of regularizing, e.g., with a thermal density operator to focus on a narrow energy window is unsuitable for our purposes, because it would require showing properties of this external thermal state that are unlikely to be accessible (even for a rigorously justified regularization strategy in Ref. [28]). To illustrate the problem, let us say that $\hat{A}=\hat{a}_{1}\otimes\hat{\mathds{1}}_{(N-1)}$ is a single-particle observable $\hat{a}_{1}$ on particle $1$. One strategy for restricting to some narrow range of energies, often useful in rigorous statements on thermalization [28, 58], is to initialize the remaining (N1)(N-1) particles in some state $\hat{\rho}\propto\hat{\mathds{1}}_{1}\otimes\hat{\rho}_{N-1}$ with narrow energy support and measure the autocorrelator of $\widehat{\delta a}_{1}=\hat{a}_{1}-a_{\mathrm{th}}\hat{\mathds{1}}_{1}$ in this state. In attempting to derive expressions analogous to Eq. (15), we get:
@@
where $A_{\mathrm{th}}(\mathcal{E})$ is the thermal value in an energy shell of range $\mathcal{E}$ (assumed constant within the energy shell). Then we specifically show that a class of interference measures, which we generically call “quantum dynamical echoes”, related to Loschmidt echoes [59, 60, 61],
@@
can be time-averaged with completely positive weight functions $w_{+}(t)$ and $v_{+}(t)$ for linear combinations of $t_{1}$ and $t_{2}$ to constrain energy-band thermalization within the energy shell spanning $\mathcal{E}$, with any desired energy-band width $\Delta E$. This rigorously connects the properties of the energy eigenstates to certain Loschmidt echoes, which has so far remained one of the missing links between two key objects of interest that fall under the umbrella of “quantum chaos” [62], which we take to refer to a collection of loosely related notions with clear physical distinctions that are often used to study complex quantum systems (see, e.g., [63] for a summary of some links). We also show that the microscopic properties of energy levels may once again be bypassed completely, by directly relating the decay of these echoes to the thermalization of any physical basis states within the energy shell $\mathcal{E}$, as described for time-averaged thermalization in Summary 6.3, with its generalization to thermal equilibrium in Sec. 7.
@@
This strategy can also account for macroscopic conserved charges [^5] As all Hamiltonians conserve linear combinations of the energy projectors $\lvert E_{n}\rangle\langle E_{n}\rvert$ (including all projectors in degenerate eigenspaces), by a conserved charge we strictly just mean some observable $\hat{Q}$ formed by such a linear combination — which commutes with the Hamiltonian $\hat{H},\hat{Q}]=0$ — on which the thermal value of our observable $\hat{A}$ of interest may strongly depend. By “macroscopic”, we mean a conserved charge with an inaccessibly large number of eigenvalues. For discrete conserved charges (with an accessible number of eigenvalues), it should be straightforward to project onto a desired eigenspace of the charge.: in the presence of a single conserved charge $\hat{Q}$ (for example) such that the thermal value $A_{\mathrm{th}}(\mathcal{E},\mathcal{Q})$ depends [17, 64] on $Q$, being effectively constant over a range of eigenvalues $Q\in\mathcal{Q}$, we can update Eq. (20) to the form:
@@
Here, it is not essential for the charge $\hat{Q}_{\text{exp}}$ that may be accessible in experiments to be the exact conserved charge $\hat{Q}$, but just for the transformation $\exp(-i\hat{Q}_{\text{exp}}s)$ generated by it for accessible values of $s$ to be sufficiently close to the true symmetry transformation $\exp(-i\hat{Q}s)$ to allow a determination of the above echo to within experimental errors (see also [65]).
In summary, we develop a conceptual framework for quantum statistical mechanics that relies entirely on the measurable dynamical properties of a single state (per observable), as illustrated through explicit measurement protocols. This parallels what we find to be the simplest dynamical justification for classical statistical mechanics [8] (Eq. (1)), but with additional interesting consequences of interference effects in the quantum case that allow us to explicitly address finite time intervals, a finite energy resolution, and the presence of accessible conserved charges in an exact manner.
@@ 1.3 Organization of this paper
Sec. 3 reviews the conventional quantum statement of eigenstate thermalization with infinite time averages, but also our version (Definition 3.2) of a stronger statement for thermalization in degenerate eigenspaces that seems not to be widely known at all, whose variants have appeared before in Refs. [41, 50, 66]. Despite its relevance only for infinite timescales, we will use this notion to motivate energy-band thermalization by regarding energy bands as “approximate” degenerate eigenspaces over small timescales that are not sensitive to their resolution.
@@
For the Appendices, Appendix A derives a finite-time Mazur-Suzuki inequality for a finite dimensional quantum system, rigorously bounding completely positive time averages of autocorrelators in terms of approximately conserved charges. Appendix B illustrates an application of our results to show the thermalization of local observables in almost all (physical) initial states of dual-unitary quantum circuits. Appendix C analyzes the connection between energy-band thermalization and eigenspace thermalization, and concludes that eigenspace thermalization is not always accessible even given energy-band thermalization; consequently, attempting to bypass the energy levels and eigenstates entirely may be in our best interest for an experimentally accessible formulation of quantum statistical mechanics. Finally, as all our proofs are relatively straightforward (and repeated) applications of the triangle and Cauchy-Schwarz inequalities [67], they are relegated to Appendix D, although what we perceive to be nontrivial physical statements are derived or explained in the course of the main text or the relevant Appendices.
@@ 2 Classical “eigenstate” thermalization without ergodicity
Now, with a primary view of motivating our quantum results, we will describe one possible approach to the classical autocorrelator, Eq. (1), as the determiner of thermalization in classical dynamical systems. Rather than take a more direct route [8], we will follow a somewhat contrived path by first considering a classical notion of eigenstate thermalization as a possible explanation for thermalization. While classical eigenstate thermalization has been discussed before in, e.g., Ref. [50] assuming the ergodic hypothesis, it is crucial for our purposes to formulate this notion for more general non-ergodic systems to resemble the quantum situation and demonstrate independence from ergodicity.
@@ 2.1 Setup: classical ergodicity
Figure 2: Schematic depiction of thermalization without ergodicity in a classical phase space $\mathcal{P}$ with 55 ergodic subsets $\mathcal{P}_{k}$. The observable $\Pi_{A}$ is considered to be equally distributed over these ergodic subsets (i.e., shows eigenstate thermalization), and must therefore thermalize in almost all initial states. The observable $\Pi_{B}$ is restricted to a few of these subsets, and cannot thermalize almost everywhere. Separately, this figure also provides schematic intuition for the Mazur-Suzuki inequality [Eq. (18) and App. A], where the conserved quantities are regarded to be projectors onto the ergodic subspaces $Q_{k}=\Pi_{\mathcal{P}_{k}}$, among which we have access to, say, $k=3,4,5$. Here, $\Pi_{B}$ has a large overlap with these conserved quantities and its autocorrelator may fail to decay to its thermal value by the inequality, while $\Pi_{A}$ has sufficiently low overlap with all these conserved quantities to still allow autocorrelator thermalization (which in turn would imply eigenstate thermalization as well as thermalization).
Ergodicity is a difficult property to either prove or verify in any sufficiently large phase space, not least because of the sensitivity required to ensure that every infinitesimal region of $\mathcal{P}$ is visited by generic trajectories (in a 33D gas of $N\to\infty$ particles, this would require having precise knowledge of the $3N$ positions and momenta of every particle). Instead of relying on ergodicity, let us ask when Eq. (29) can be true for a given measurement $A$ in a non-ergodic system. This is of extreme physical relevance: all nontrivial Hamiltonian systems are non-ergodic due to the conservation of energy, as well as the frequent presence of various other conservation laws (such as total momentum or angular momentum, depending on the system under consideration and nature of interactions). Further, experimental uncertainties ensure that any initial state is statistically distributed over a range of energies $\Delta E$. Moreover, as ergodicity in the dynamical sense is difficult (in most cases, near-impossible) to establish for a sufficiently complex system, so a given system may as well be nonergodic for all practical purposes. In this case, the phase space can be decomposed into an unknown number of subsets $\mathcal{P}_{k}$:
@@
each of which is ergodic. These are typically surfaces of fixed energy, as well as fixed values of other conserved quantities (if present). Through a slight misuse of terminology that anticipates quantum mechanics, we will call the sets $\mathcal{P}_{k}$ — or more precisely, uniform distributions supported entirely on one of these sets — the “energy eigenstates” of the classical Hamiltonian (for example, because they correspond to the classical limit of the quantum energy eigenstates, and satisfy similar properties such as having a definite value of energy and being invariant under time evolution [^6] This parallel can be taken further in the Koopman-von Neumann framework [5] of classical Hilbert spaces comprised of functions on $\mathcal{P}$, in which classical dynamics has well-defined eigenstates.). An identification between constant energy surfaces and eigenstates was proposed in Ref. [50] in a related context, assuming ergodicity on each energy surface; for our perspective, it is instead crucial that we identify eigenstates with the subsets $\mathcal{P}_{k}$ especially if the system is non-ergodic.
For simplicity, we will assume that there is a finite number $M$ of $\mathcal{P}_{k}$, each with nonzero measure [^7] Any subsets of zero measure can be trivially absorbed into any nonzero measure subset without affecting its ergodicity. $\mu(\mathcal{P}_{k})>0$; formally, this simplifies the technical details of several arguments, and the case of a continuum of energy surfaces (typical in classical systems) can be recovered via a continuum limit $M\to\infty$ without altering any of our conclusions. This assumption allows us to focus on the physics of the problem, rather than the formal definition of induced measures on subsets and their integration. It also further anticipates quantum mechanics, in which there is a minimum phase space volume [68, 69] connected to the inverse purity of a pure state. With this simplification, due to the ergodicity of each subset, any observable $\Pi_{A}$ satisfies
@@
If Eq. (34) is satisfied, we say that the observable $\Pi_{A}$ shows eigenstate thermalization: its expectation value in every ergodic subset $\mathcal{P}_{k}$ is the thermal value $\mu(A)/\mu(\mathcal{P})$. Though classically such a notion is not particularly useful, because it tends to be the case that a typical accessible observable $A$ attains its thermal value in almost all states $x\in\mathcal{P}$ in the phase space (and therefore trivially for the energy eigenstates) [8, 9], we will continue to formally entertain this notion in anticipation of later sections.
@@
See App. D.1.2. The intuition behind the proof (which is reminiscent of a proof of semiclassical wavefunction ergodicity theorems [40, 41, 46], which we will revisit for quantum thermalization) is as follows. The difference between the left and right hand sides of Eq. (35),